OpenStudy (ilovecake):

What is 4/9 divided by 1/3 A.4/27 B.1 2/3 C.1 1/3 D.7/9

OpenStudy (igreen):

It's the same as multiplying by the reciprocal of 1/3, which is 3. \(\dfrac{4}{9} \times 3 =~?\)

OpenStudy (ilovecake):

I think the answer might be C.

OpenStudy (igreen):

Yep, C is correct.

OpenStudy (ilovecake):

ok thanks!

OpenStudy (anonymous):

hey @iGreen you wanna help me with something?

OpenStudy (igreen):

\(\dfrac{4}{9} \times 3 \rightarrow \dfrac{12}{9} \rightarrow 1\dfrac{1}{3}\)

OpenStudy (ilovecake):

thanks

OpenStudy (igreen):

I might be able to, tag me in your own question. @breathecarolina